### v1

# Osmotic forces modify lipid membrane fluctuations

Abstract. In hydrodynamic descriptions of lipid bilayers, the membrane is often approximated as being impermeable to the surrounding, solute-containing fluid. However, biological and in vitro lipid membranes are influenced by their permeability and the resultant osmotic forces—whose effects remain poorly understood. Here, we study the dynamics of a fluctuating, planar lipid membrane that is ideally selective: fluid can pass through it, while the electrically-neutral solutes cannot. We find that the canonical membrane relaxation mode, in which internal membrane forces are balanced by fluid drag, no longer exists over all wavenumbers. Rather, this mode only exists when it is slower than solute diffusion—corresponding to a finite range of wavenumbers. The well-known equipartition result quantifying the size of membrane undulations due to thermal perturbations is consequently limited in its validity to the aforementioned range. Moreover, this range shrinks as the membrane surface tension is increased, and above a critical tension the membrane mode vanishes. Our findings are relevant when interpreting experimental measurements of membrane fluctuations, especially in vesicles at moderate to high tensions.
[…]
Introduction
Biological membranes are unique materials that surround the cell, and compartmentalize its internal organelles. While such membranes are slowly permeable to water [1], large neutral molecules (such as sugars) cannot easily pass through them [2]. The semipermeability of lipid membranes is relevant to both biological and in vitro scenarios—as in both cases, membranes are essentially always surrounded by solutes. While the general equations governing the coupled membrane, fluid, and solute dynamics were recently obtained [3], there is not yet a systematic characterization of membrane behavior. As a result, the effects of permeability and osmosis on lipid membrane dynamics remain poorly understood.
Significance
Lipid membranes are ubiquitous in biological and in vitro systems. Though they are often approximated as impermeable, water can slowly pass through the membrane. In this work, we study the consequences of lipid membranes being permeable to water, but not solutes (such as sugars). We find that when solutes are fast relative to the membrane—which occurs at low surface tensions and intermediate wavelengths—membrane undulations are indistinguishable from the impermeable prediction. However, at large and small wavelengths or high tensions, solutes are slow relative to the membrane and the impermeable result does not hold. Our findings are relevant to the many experimental investigations that apply impermeable results across all wavelengths, and all tensions, in their characterization.
The present study investigates how planar lipid membranes act on, and are influenced by, their surrounding environment. Flat membranes are found in various biological settings, including the endoplasmic reticulum [4], Golgi complex [5], and neuromuscular junction [6]. In addition, theories of impermeable bilayer fluctuations about a flat surface are often used to describe the thermal undulations of giant unilamellar vesicles (GUVs). We begin by investigating an impermeable membrane, where the dynamics of the solute are not considered. By starting with a well-understood system (see e.g. Refs. [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]), we are better equipped to study our primary system of interest: the semipermeable membrane. The temporal evolution of the normal modes is solved for within the linear theory, and the relative time scale of membrane and solute relaxation is found to dictate the nature of the coupled dynamics. When solutes relax more quickly, the dynamics are indistinguishable from the impermeable scenario. It is then straightforward to present an effective Langevin equation including thermal perturbations, and obtain the magnitude of the resultant height fluctuations. In contrast, when solutes are slow relative to the membrane, the impermeable dynamics are unphysical and the resultant behavior is outside the scope of a normal mode analysis within a linear description. As a result, the character of thermal undulations—long assumed to be that of an impermeable membrane when analyzing fluctuating GUVs—is no longer known. Experimental consequences of our findings are highlighted, and avenues for future work are discussed.
The impermeable membrane
Consider a nearly planar lipid membrane surrounded by water. For the present analysis, we neglect solutes and make the standard approximation that the membrane is impermeable to fluid. Water is characterized by its three-dimensional mass density $\rho_{\mkern-1.0mu\text{f}}$, shear viscosity $\mu_{\mkern-1.0mu\text{f}}$, and kinematic viscosity $\nu_{\mkern-1.0mu\text{f}}:=\mu_{\mkern-1.0mu\text{f}}/\rho_{\mkern-1.0mu\text{f}}$. Relevant membrane parameters are the patch length $\ell_{\text{c}}$, two-dimensional (2D) areal density $\rho_{\mkern-1.0mu\text{m}}$, surface tension $\lambda_{\text{c}}$, and bending modulus $k_{\text{b}}$ [20]. It is important to note that the surface tension is not a membrane property, but rather takes the requisite value to enforce areal incompressibility [21, 22]. As a result, the membrane surface tension can span a wide range of values, and is often tuned in experiments [23, 24]. Characteristic values of all parameters are provided in Table 1.
Our base state is an unperturbed membrane sheet lying in the xxyy plane, surrounded on both sides by a semi-infinite stationary fluid. We seek to understand how the coupled membrane and fluid system responds to small perturbations. To this end, we determine the linearized dynamics of the coupled membrane and fluid system about the base configuration. In what follows, an abbreviated analysis of the governing equations is presented; see §2 of the Supplemental Material (SM) [25] for further details.
The governing equations
The position $\bm{x}$ of the perturbed membrane is given by $\bm{x}(x,y,t)=x\mkern 1.0mu\bm{e}_{x}+y\mkern 1.0mu\bm{e}_{y}+h(x,y,t)\mkern 1.0mu\bm{e}_{z}$, where $x,y\in[0,\ell_{\text{c}}]$ and $\lvert h\rvert\ll\ell_{\text{c}}$ by construction. For planar geometries with a stationary base state, the in-plane and out-of-plane membrane equations are decoupled [26, 27]. To solve for the time evolution of the membrane height, we require only the $z$-component of the linear momentum balance of the membrane—often called the shape equation and given by [27]
$\rho_{\mkern-1.0mu\text{m}}\mkern 1.0muh_{\mkern-1.0mu,\mkern 1.0mutt}\,=\,\lambda_{\text{c}}\mkern 1.0muh_{\mkern-1.0mu,\mkern 1.0mu\alpha\alpha}\,-\,\dfrac{k_{\text{b}}}{2}\mkern 1.0muh_{\mkern-1.0mu,\mkern 1.0mu\alpha\alpha\beta\beta}\,-\,\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket~.$ (1)
In Eq. (1), $\alpha,\beta\in{1,2}$ denote directions in the xxyy plane, with repeated indices summed over such that (for example) $h_{\mkern-1.0mu,\mkern 1.0mu\alpha\alpha}=h_{\mkern-1.0mu,\mkern 1.0muxx}+h_{\mkern-1.0mu,\mkern 1.0muyy}$. Here $(\,\bm{\cdot}\,)_{,t}:=\partial(\,\bm{\cdot}\,)/\partial t$ is the partial time derivative and $\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket:=p^{+}-p^{-}$ is the difference in the fluid pressure above (++) and below (-) the membrane. The pressure and velocity fields, $p^{\pm}$ and $v^{\pm}_{\mkern-1.0muj}$ (where Roman indices span $\{1,2,3\}$ and repeated indices are summed over), satisfy the incompressible Navier–Stokes equations linearized about the base state: $v^{\pm}_{\mkern-1.0muj,\mkern 1.0muj}=0$ and $\rho_{\mkern-1.0mu\text{f}}\mkern 1.0muv^{\pm}_{\mkern-1.0muj,\mkern 1.0mut}=\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muv^{\pm}_{\mkern-1.0muj,\mkern 1.0mukk}-p^{\pm}_{\mkern-1.0mu,\mkern 1.0muj}$. With the no-slip boundary condition $h_{\mkern-1.0mu,\mkern 1.0mut}=v^{\pm}_{\mkern-1.0muz}$ at the membrane surface, water cannot flow through the membrane and we have a well-posed set of equations for the coupled system [25, §2].
At this point, we follow standard approaches to solve for the dynamics of the system. The membrane height is decomposed into planar Fourier modes as
$h(x,y,t)\,=\,\sum_{\bm{q}}\hat{h}_{\bm{q}}\,e^{i\mkern 1.0mu(q_{x}x\,+\,q_{y}y)}\,e^{-\omega_{\bm{q}}t}~,$ (2)
where the wavevector $\bm{q}=(q_{x},q_{y})$ has corresponding wavenumber $q:=(q_{x}^{\,2}+q_{y}^{\,2})^{1/2}$, $\omega_{\bm{q}}$ is the associated relaxation frequency, and normal modes are assumed independent. Fluid unknowns are also decomposed into planar Fourier modes, with the $z$-dependence to be solved for. The fluid pressure, for example, is expressed as
$p^{\pm}(x,y,z,t)\,=\,\sum_{\bm{q}}\hat{p}^{\mkern 1.0mu\pm}_{\bm{q}}(z)\,e^{i\mkern 1.0mu(q_{x}x\,+\,q_{y}y)}\,e^{-\omega_{\bm{q}}t}~.$ (3)
Our main task is to solve for all frequencies $\omega_{\bm{q}}$ satisfying the governing equations. In doing so, we also solve for the pressure and velocity fields in the fluid.
Table 1: Membrane, fluid, and solute parameters.
Parameter Sym. Value Ref.
Patch size $\ell_{\text{c}}$ $10^{2}\text{--}10^{3}$ nm [4, 6]
Density $\rho_{\mkern-1.0mu\text{m}}$ $10^{-8}$ pg/nm2 [28]
Bending modulus $k_{\text{b}}$ $10^{2}~\text{pN}\!\cdot\!\text{nm}$ [29]
Surface tension $\lambda_{\text{c}}$ $10^{-4}$$10^{-1}$ pN/nm [29, 30]
[…]
Density $\rho_{\mkern-1.0mu\text{f}}$ $10^{-9}$ pg/nm3
[…]
Kinematic viscosity $\nu_{\mkern-1.0mu\text{f}}$ $10^{6}$ nm2/µsec
Concentration $c_{0}$ $10^{-1}$ nm-3 [32]
Diffusivity $D$ $10^{3}$ nm2/µsec [33]
The frequency solutions
As we limit ourselves to the linearized dynamics, all unknowns of a particular wavevector $\bm{q}$ are proportional to the membrane height $\hat{h}_{\bm{q}}$. We assume $\hat{h}_{\bm{q}}\neq 0$ and combine all equations governing membrane and fluid into a single equation for $\omega_{\bm{q}}$, namely the dispersion relation (see §2.3 of the SM [25]). Given the physically-motivated requirement that all fluid perturbations decay far from the membrane surface, there are two frequencies at each wavenumber, which are plotted in Fig. 1. The aforementioned equation for $\omega_{\bm{q}}$ is not straightforward to interpret—though we find the dynamics are well-approximated by
$\rho_{\mkern-1.0mu\text{eff}}\,\omega_{\bm{q}}^{\,2}\,-\,4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq\,\omega_{\bm{q}}\,+\,E\,=\,0~.$ (4)
Equation (4), with the temporal ansatz $h\sim e^{-\omega t}$ in Eq. (2), reveals the dynamics of the membrane are analogous to that of the spring–mass–damper from introductory mechanics. Here $\rho_{\mkern-1.0mu\text{eff}}:=\rho_{\mkern-1.0mu\text{m}}+\mkern 1.0mu4\mkern 1.0mu\rho_{\mkern-1.0mu\text{f}}/q$ approximates the combined inertia of membrane and fluid, $4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq$ is the well-known hydrodynamic drag from the surrounding fluid, and $E:=\lambda_{\text{c}}\mkern 1.0muq^{2}+\tfrac{1}{2}\mkern 1.0muk_{\text{b}}\mkern 1.0muq^{4}$ captures the energetics of small membrane fluctuations: $\tfrac{1}{2}\,\ell_{\text{c}}^{\,2}\mkern 1.0muE\mkern 1.0mu\lvert\hat{h}_{\bm{q}}^{\,2}\rvert$ is the energy of mode $\bm{q}$ to quadratic order [34]. Since inertial effects are small, Eq. (4) has the two approximate solutions
$\tilde{\omega}_{\mkern-2.0mu\rho}\,:=\,\dfrac{4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq}{\rho_{\mkern-1.0mu\text{eff}}}\qquad\text{and}\qquad\tilde{\omega}_{\mkern-1.0mu\text{m}}\,:=\,\dfrac{E}{4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq}~.$ (5)
In Eq. (5), the ‘tilde’ accent indicates that Eq. (4) is only an approximation of the true dynamics; the subscripts $\rho$ and $\mathrm{m}$ designate the inertial and membrane branches. As shown in Fig. 1, the frequencies in Eq. (5) are a good approximation to the true solution.
[…]
The membrane response. Following Eqs. (4) and (5), let us denote the two exact frequency branches as $\omega_{\mkern-2.0mu\rho}$ and $\omega_{\mkern-1.0mu\text{m}}$. The membrane height, in real space, is then given by [cf. Eq. (2)]
$h(x,y,t)\,=\,\sum_{\bm{q}}\!\Big(\hat{h}_{\bm{q}}^{\rho}\mkern 1.0mue^{-\omega_{\mkern-2.0mu\rho}t}\mkern 1.0mu+\mkern 1.0mu\hat{h}_{\bm{q}}^{\text{m}}\mkern 1.0mue^{-\omega_{\mkern-1.0mu\text{m}}t}\Big)\mkern 1.0mue^{i\mkern 1.0mu(q_{x}x\,+\,q_{y}y)},$ (6)
where $\hat{h}_{\bm{q}}^{\rho}$ and $\hat{h}_{\bm{q}}^{\text{m}}$ are determined upon specification of the initial conditions: $h$ and $\partial h/\partial t$ at time $t=0$. Since $\omega_{\mkern-2.0mu\rho}\gg\omega_{\mkern-1.0mu\text{m}}$ over all wavenumbers, the inertial branch quickly decays and the dynamics of the coupled system are captured by the membrane branch. Prior studies are accordingly justified in neglecting inertial effects throughout their investigations of planar lipid membrane behavior [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 35, 36, 37, 38, 39].
The fluid response. To understand the dynamics of the fluid as the membrane relaxes, consider two disturbances to a specific mode $\bm{q}$. In the first case, $\hat{h}^{\mathrm{m}}_{\bm{q}}=0$ and only the inertial mode is excited, while in the second case $\hat{h}^{\rho}_{\bm{q}}=0$ and only the membrane mode is excited [see Eq. (6)]. In both cases, the pressure field (3) is given by $\hat{p}^{\mkern 1.0mu\pm}_{\bm{q}}=\mp\tfrac{1}{2}(\rho_{\mkern-1.0mu\text{m}}\mkern 1.0mu\omega_{\bm{q}}^{2}+E)\,\hat{h}_{\bm{q}}\mkern 1.0mue^{-q\lvert z\rvert}$. The magnitude of the pressure at $z=0$ determined by Eq. (1), and the decay length $1/q$ is a consequence of the pressure field being harmonic: $p^{\pm}_{\mkern-1.0mu,\mkern 1.0mujj}=0$. In contrast, fluid velocities decay over the two length scales $1/q$ and $1/k_{\mkern-1.0mu\nu}$, where $k_{\mkern-1.0mu\nu}:=(q^{2}-\omega/\nu_{\mkern-1.0mu\text{f}})^{1/2}$ is a modified wavenumber that accounts for the inertia of the fluid (see §94 of Ref. [40] and §2.1 of the SM [25]). As a consequence of the time scale separation $\omega_{\mkern-2.0mu\rho}\gg\omega_{\mkern-1.0mu\text{m}}$, $1/\bar{q}\approx 1/k_{\mkern-1.0mu\nu}$ on the membrane branch and $1/\bar{q}\ll 1/k_{\mkern-1.0mu\nu}$ on the inertial branch.
(a) impermeable overlay
(b) dome comparing frequencies
(c) semipermeable branches
Figure 2: Relevance of the diffusive time scale to the dynamics of the semipermeable system. (a) An overlay of $\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ (solid green line) on top of the impermeable frequencies in Fig. 1. There are two wavenumbers for which the impermeable membrane frequency $\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}$ equals $\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$; approximating the former as $\tilde{\omega}_{\mkern-1.0mu\text{m}}$ yields the two crossover wavenumbers $q_{0}^{-}$ and $q_{0}^{+}$ (10)—shown as vertical, dashed lines. (b) Plot of $q_{0}^{\pm}$ over a range of the base membrane tension $\lambda_{\text{c}}$, which varies across systems. The surface tension from (a) of $\lambda_{\text{c}}=2\!\cdot\!10^{-3}$ pN//nm is shown as the horizontal dashed line. There exists a critical tension $\lambda_{\text{c}}^{\mkern-1.0mu\ast}$ (11) above which $\tilde{\omega}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ for all wavenumbers. (c) Plot of the real parts of the semipermeable frequencies (dotted or dashed solid lines), overlaid on top of the impermeable frequencies (thick, transparent bands) for comparison. The inertial branch is essentially unchanged. The membrane branch only exists when $\omega_{\mkern-1.0mu\text{m}}<\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$—and is thus confined to the range $q\in(q_{0}^{-},q_{0}^{+})$. For base membrane tensions $\lambda_{\text{c}}>\lambda_{\text{c}}^{\mkern-1.0mu\ast}$, the membrane branch vanishes altogether.
The semipermeable membrane
[…]
The governing equations
The membrane shape equation (1) is unchanged in the presence of solutes [3]. The perturbed concentrations $c^{\pm}$ above and below the membrane, with $\lvert c^{\pm}\rvert\ll c_{0}$, evolve according to the diffusion equation $c^{\pm}_{\mkern-1.0mu,\mkern 1.0mut}=D\mkern 1.0muc^{\pm}_{\mkern-1.0mu,\mkern 1.0mujj}$. Solute impermeability is enforced by requiring the total solute flux through the membrane—comprised of diffusive and convective contributions—to be zero:
$D\mkern 1.0muc^{\pm}_{\mkern-1.0mu,\mkern 1.0muz}\,+\,c_{0}\mkern 1.0mu\big(v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut}\big)\,=\,0~.$ (7)
In Eq. (7), velocities and concentration gradients are evaluated at the membrane surface. The equation governing water permeation through an arbitrarily curved and deforming lipid membrane was recently obtained by A.M. Alkadri and K.K. Mandadapu [3], who extended several seminal findings [41, 42, 43]. These authors used techniques from irreversible thermodynamics and differential geometry to show fluid flow through the membrane is driven by differences in hydrodynamic tractions and osmotic pressures across the membrane surface. For nearly planar geometries and an ideally selective membrane, the findings from Ref. [3] simplify to
$\kappa\mkern 1.0mu\big(v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut}\big)\,=\,k_{\text{B}}\theta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket\,-\,\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket~,$ (8)
where $k_{\text{B}}$ is Boltzmann’s constant, $\theta$ is the absolute temperature, and $\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket:=c^{+}-c^{-}$ is the jump in solute concentration at the membrane surface. In the language of irreversible thermodynamics [44, 45], Eq. (8) presents a linear relationship between the thermodynamic driving force $(k_{\text{B}}\theta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket-\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket)$ and the corresponding thermodynamic flux $(v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut})$, with the latter a measure of fluid flow through the membrane. The phenomenological parameter $\kappa$, understood as a transport coefficient [3], captures the resistance to flow: for a given driving force, a larger value of $\kappa$ corresponds to a smaller water flux through the membrane. Assuming the usual Stokes flow scaling of $\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq^{2}v\sim q\mkern 1.0mup$, in the limit of $\kappa\rightarrow\infty$ or $\mathcal{P}\rightarrow 0$ we recover the impermeable coupling condition $v^{\pm}_{\mkern-1.0muz}=h_{\mkern-1.0mu,\mkern 1.0mut}$.
The diffusive time scale
The presence of solutes in the surrounding fluid introduces a diffusive frequency
$\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}\,:=\,q^{2}D~.$ (9)
Figure 2(a) shows $\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ plotted on top of the curves in Fig. 1; note the diffusive and impermeable membrane frequencies can be comparable. Let us denote $q_{0}^{\pm}$ as the two wavenumbers at which the diffusive and membrane time scales are equal. From Eqs. (5) and (9), we find
$q_{0}^{\pm}\,=\,\dfrac{4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muD}{k_{\text{b}}}\,\bigg[\mkern 1.0mu1\,\pm\,\bigg(1\,-\,\dfrac{\lambda_{\text{c}}\mkern 1.0muk_{\text{b}}}{8\mkern 1.0mu{\mu_{\mkern-1.0mu\text{f}}}^{\!2}D^{2}}\bigg)^{\!\!1/2\,\,}\mkern 1.0mu\bigg]~,$ (10)
with $\tilde{\omega}_{\mkern-1.0mu\text{m}}<\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ when $q\in(q_{0}^{-},q_{0}^{+})$, and $\tilde{\omega}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ otherwise. Since the base surface tension $\lambda_{\text{c}}$ is not a material parameter, but rather varies—and can be tuned—across systems, $q_{0}^{\pm}$ is plotted at different tensions $\lambda_{\text{c}}$ in Fig. 2(b) to reveal a dome-like envelope. There exists a critical base tension
$\lambda_{\text{c}}^{\mkern-1.0mu\ast}\,:=\,\dfrac{8\mkern 1.0mu{\mu_{\mkern-1.0mu\text{f}}}^{\!2}D^{2}}{k_{\text{b}}}\,\sim\,6\cdot 10^{-2}\text{ pN/nm}$ (11)
at the top of the dome, above which no real $q_{0}$ exists. When the base membrane tension $\lambda_{\text{c}}$ is larger than $\lambda_{\text{c}}^{\mkern-1.0mu\ast}$, $\tilde{\omega}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ over all wavenumbers. As we will see, the relative values of $\tilde{\omega}_{\mkern-1.0mu\text{m}}$ and $\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$—and thus whether a mode is “inside” (i.e. below) or “outside” (i.e. above) the dome in Fig. 2(b)—govern different regimes of membrane behavior.
The frequency solutions
Membrane and fluid unknowns are expressed as in the impermeable analysis, with the concentration expanded in planar normal modes as [cf. Eq. (3)]
$c^{\pm}(x,y,z,t)\,=\,\sum_{\bm{q}}\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}(z)\,e^{i\mkern 1.0mu(q_{x}x\,+\,q_{y}y)}\,e^{-\omega_{\bm{q}}t}~.$ (12)
In our linearized description, all perturbed bulk quantities are again proportional to the membrane height. Assuming a nontrivial solution ($\hat{h}_{\bm{q}}\neq 0$), it is straightforward to arrive at the dispersion relation, as shown in §3.3 of the SM [25]. While $\omega_{\bm{q}}$ cannot be solved for analytically, we take advantage of the smallness of $\mathcal{P}$: since the membrane is nearly impermeable, semipermeable frequencies are close to their impermeable counterparts. In what follows, we investigate how the dynamics of the impermeable branches are altered by the presence of solutes in the surrounding medium.
The inertial branch. We begin by assuming concentration modes evolve with the impermeable inertial frequency $\omega^{\text{imp}}_{\mkern-2.0mu\rho}$. The diffusion equation for mode $\bm{q}$ then simplifies to
$\dfrac{\text{d}^{2}\mkern 1.0mu{\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}}}{\text{d}\mkern 1.0mu{z}^{2}}\,=\,-q^{2}\mkern 1.0mu\tilde{\eta}^{2}\mkern 1.0mu\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}~,\qquad\text{where}\qquad\tilde{\eta}\,:=\mkern 1.0mu\bigg(\dfrac{\omega^{\text{imp}}_{\mkern-2.0mu\rho}}{\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}}\,-\,1\bigg)^{\!\!1\mkern-1.0mu/\mkern-1.0mu2}$ (13)
is a positive, real, dimensionless parameter defined for notational convenience. Equation (13) admits the two purely oscillatory solutions $\hat{c}^{\mkern 1.0mu+}_{\bm{q}}\sim e^{\pm i\mkern 1.0muq\mkern 1.0mu\tilde{\eta}\mkern 1.0muz}$. As these standing waves do not decay as $z\rightarrow\infty$, we seek the lowest-order concentration correction due to the semipermeability of the membrane.***For conciseness, only the concentration field above the membrane is presented; see the SM [25, §§ 3.2–3.3] for all details.
We first recognize that water permeation alters the hydrodynamic drag on the membrane by the surrounding fluid. Equations (7) and (8) together yield $\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket=k_{\text{B}}\theta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket-(\kappa\mkern 1.0muD/c_{0})\mkern 1.0muc^{\pm}_{\mkern-1.0mu,\mkern 1.0muz}\mkern 1.0mu$, which—along with Eq. (1)—shows standing waves in the concentration add a hydrodynamic drag force 90 out of phase with the membrane velocity. We show in the SM [25, §3.3.2] that an approximate dispersion relation for the semipermeable membrane is given by [cf. Eq. (4)]
$\rho_{\mkern-1.0mu\text{eff}}\,\omega_{\bm{q}}^{\,2}\,-\,4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq\,\omega_{\bm{q}}\,+\,E\,=\,\pm\,\dfrac{8\mkern 1.0mui\mkern 1.0mu\mathcal{P}\mkern 1.0mu\mathcal{S}}{\tilde{\eta}}\,(4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq)\,\omega_{\bm{q}}~.$ (14)
In Eq. (14), $\mathcal{S}:=k_{\text{B}}\theta\mkern 1.0muc_{0}/(D\mkern 1.0mu\kappa\mkern 1.0muq)\ll 1$ as a dimensionless parameter that arises when Eqs. (7) and (8) are combined. The solution of Eq. (14) closest to $\omega^{\text{imp}}_{\mkern-2.0mu\rho}$, to lowest order in the product $\mathcal{P}\mkern 1.0mu\mathcal{S}$ of small parameters, is given by
[…]
Membrane permeability and osmotic forces accordingly introduce exceedingly slow temporal oscillations to the inertial dynamics of the coupled system.
The modification to the inertial frequency in Eq. (15) also alters the spatial concentration of the solutes. Substituting Eq. (15) into the diffusion equation and solving for the concentration field yields
$\hat{c}^{\mkern 1.0mu+}_{\bm{q}}\,$ $\sim\,\exp\mkern 1.0mu\bigg\{\pm\,i\mkern 1.0muq\mkern 1.0mu\tilde{\eta}\mkern 1.0muz\,-\,\dfrac{4\mkern 1.0mu\mathcal{P}\mkern 1.0mu\mathcal{S}\mkern 1.0muq\mkern 1.0muz}{1-\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}/\omega^{\text{imp}}_{\mkern-2.0mu\rho}}\bigg\}~.$ (16)
Importantly, $\hat{c}^{\mkern 1.0mu+}_{\bm{q}}\rightarrow 0$ as $z\rightarrow\infty$ and concentration perturbations due to the fluctuating membrane decay far from the membrane surface. We have thus found a physically meaningful solution for both the frequency (15) and concentration (16) over all wavenumbers. The validity of our approximate analysis is confirmed in Fig. 2(c): the impermeable branch is visually indistinguishable from the real portion of its semipermeable counterpart.
The membrane branch when diffusion is slow. As in the analysis of the inertial branch, we start by assuming the solute concentration evolves at the frequency of the impermeable system—in this case, the membrane frequency $\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}$. For wavenumbers $q\notin(q_{0}^{-},q_{0}^{+})$, diffusion is slow relative to the membrane and $\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$. The solute diffusion equation once again simplifies to Eq. (13), where now $\tilde{\eta}$ involves $\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}$ instead of $\omega^{\text{imp}}_{\mkern-2.0mu\rho}$. With this change, and following a similar analysis as that described in the inertial scenario, we once again arrive at Eq. (14). The membrane frequency is then found to be given by
[…]
Importantly, the sign of the imaginary part of $\omega_{\mkern-1.0mu\text{m}}$ is opposite that of $\omega_{\mkern-2.0mu\rho}$ [cf. Eq. (15)], and the out-of-phase drag force alters the inertial and membrane dynamics in different ways. Substituting Eq. (17) into the diffusion equation gives the concentration field [cf. Eq. (16)]
$\hat{c}^{\mkern 1.0mu+}_{\bm{q}}\,$ $\sim\,\exp\mkern 1.0mu\bigg\{\pm\,i\mkern 1.0muq\mkern 1.0mu\tilde{\eta}\mkern 1.0muz\,+\,\dfrac{4\mkern 1.0mu\mathcal{P}\mkern 1.0mu\mathcal{S}\mkern 1.0muq\mkern 1.0muz}{1-\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}/\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}}\bigg\}~.$ (18)
The concentration field in Eq. (18) diverges as $z\rightarrow\infty$, and is not physically meaningful. As a consequence, the membrane branch vanishes when $q\notin(q_{0}^{-},q_{0}^{+})$. A representative result at low tension is shown in Fig. 2(c), and at tensions larger than $\lambda_{\text{c}}^{\mkern-1.0mu\ast}$ (11) the membrane branch vanishes entirely.
The membrane branch when diffusion is fast. When $\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}<\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ and $q\in(q_{0}^{-},q_{0}^{+})$, the analysis of the coupled system is straightforward in the limit of small $\mathcal{P}$. The diffusion equation, assuming $\omega=\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}$ at lowest order, simplifies to $d^{2}\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}/dz^{2}=(q^{2}-\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}/D)\mkern 1.0mu\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}$. We thus find the exponentially decaying concentration solutions
$\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}\,\sim\,\exp\mkern 1.0mu\bigg\{\mp q\mkern 1.0muz\,\bigg(1\,-\,\dfrac{\omega^{\text{imp}}_{\mkern-1.0mu\text{m}}}{\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}}\,\bigg)^{\!\!1\mkern-1.0mu/\mkern-1.0mu2}\,\bigg\}~.$ (19)
With no oscillations in the lowest-order concentration field, modifications to the hydrodynamic drag felt by the membrane are in phase with the membrane velocity. Moreover, as such corrections are of order $\mathcal{P}\!\ll\!1$, we expect the membrane branch to be indistinguishable from its impermeable counterpart—as confirmed in Fig. 2(c).
Theoretical implications. As mentioned previously, solutes are present in the fluid surrounding nearly all artificial and biological membranes. Our findings demonstrate such systems can only be approximated as impermeable for wavenumbers “inside the dome,” where $q\in(q_{0}^{-},q_{0}^{+})$ and there is a slow frequency $\omega\approx E/(4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq)$. In contrast, only a quickly decaying inertial mode exists for wavenumbers $q\notin(q_{0}^{-},q_{0}^{+})$ “outside the dome.” For this case, ($i$) the membrane response is independent of $k_{\text{b}}$ and $\lambda_{\text{c}}$ at linear order, and ($i\mkern-1.0mui$) thermal perturbations (discussed subsequently) drive membrane undulations to continuously grow until saturated by nonlinear forces—both of which are surprising results. Such predictions motivate us to question whether our normal mode ansatz is indeed appropriate outside the dome. If instead the solute concentration was decomposed as $c\sim\check{c}(z,t)\mkern 1.0mue^{i\mkern 1.0mu(q_{x}x\,+\,q_{y}y)}$ [cf. Eq. (12)], then the diffusion equation requires $\partial\check{c}/\partial t=D(\partial^{2}\check{c}/\partial z^{2}-q^{2}\check{c})$. Given an infinite domain $z\in(-\infty,\infty)$, a self-similar solution is expected for $\check{c}(z,t)$—which is incompatible with our ansatz. Further analysis, likely involving nonlinear simulations of the membrane and surrounding fluid [46, 47, 48, 49, 50, 51, 52], is required to understand the dynamics of the coupled system. In any case, however, we emphasize that the standard impermeable result is not applicable for wavenumbers outside the dome.
Experimentally characterizing GUVs
In experimental investigations of lipid bilayers, one often seeks to characterize membrane material properties prior to further manipulation. When GUVs are involved, their fluctuations from thermal disturbances can be imaged and—with a theory of membrane undulations—processed to determine $k_{\text{b}}$ and $\lambda_{\text{c}}$ [29, 31, 53, 54, 55, 56, 57, 58, 59]. While most experimental investigations apply the well-understood impermeable theory over all wavenumbers, the present study reveals that only modes inside the dome should be treated as impermeable. Moreover, as the behavior of undulations outside the dome is not currently well-understood, experimental measurements involving these modes should be excluded when determining membrane properties. In what follows, we discuss how to incorporate thermal perturbations into our theory, and then analyze GUV fluctuation data from experiments.
The impermeable Langevin equation. When the lipid bilayer is treated as impermeable, its linear response to perturbations is well-known. Since inertial modes decay much more quickly than their membrane counterparts (see Eq. (6) with $\omega_{\mkern-2.0mu\rho}\gg\omega_{\mkern-1.0mu\text{m}}$), inertia can be neglected entirely; the membrane frequency $\tilde{\omega}_{\mkern-1.0mu\text{m}}$ presented in Eq. (5) is then exact. Decomposing the membrane height as $h(x,y,t)\,=\,\sum_{\bm{q}}\tilde{h}_{\bm{q}}(t)\mkern 1.0mue^{i\mkern 1.0mu(q_{x}x\,+\,q_{y}y)}$ leads to the Langevin equation [cf. Eqs. (2),(4)]
$\dfrac{\text{d}\mkern 1.0mu{\tilde{h}_{\bm{q}}}}{\text{d}\mkern 1.0mu{t}}\,+\,\tilde{\omega}_{\mkern-1.0mu\text{m}}\mkern 1.0mu\tilde{h}_{\bm{q}}\,=\,\tilde{\xi}_{\bm{q}}(t)~.$ (20)
In Eq. (20), $\tilde{\xi}_{\bm{q}}(t)$ is a complex Gaussian thermal noise satisfying the fluctuation–dissipation theorem, with two-time correlations proportional to $\delta(t-t^{\prime})\,k_{\text{B}}\theta/(4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq\,\ell_{\text{c}}^{\,2})$ [18]. Equation (20) can be solved exactly to yield [60]
$\langle\lvert\mkern 1.0mu\tilde{h}_{\bm{q}}\rvert^{2}\rangle\,=\,\dfrac{k_{\text{B}}\theta}{E\mkern 1.0mu\ell_{\text{c}}^{\,2}}~,$ (21)
which is the well-known result of the equipartition theorem as applied to planar membranes [34].
The Langevin equation inside the dome. For modes with $q\in(q_{0}^{-},q_{0}^{+})$, there is no discernible difference between the impermeable and semipermeable dynamics. Thus, Eq. (20) once again applies, and membrane fluctuations satisfy the equipartition result (21).
The Langevin equation outside the dome. As discussed previously, we are currently unsure if our normal mode ansatz is appropriate outside the dome. If the ansatz is indeed valid, then we effectively see only an inertial mode. The corresponding Langevin equation describing membrane fluctuations is given by
$\dfrac{1}{\omega^{(\text{r})}_{\mkern-2.0mu\rho}}\,\dfrac{\text{d}^{2}\mkern 1.0mu{\tilde{h}_{\bm{q}}}}{\text{d}\mkern 1.0mu{t}^{2}}\,+\,\dfrac{\text{d}\mkern 1.0mu{\tilde{h}_{\bm{q}}}}{\text{d}\mkern 1.0mu{t}}\,=\,\tilde{\xi}_{\bm{q}}(t)~,$ (22)
where $\omega^{(\text{r})}_{\mkern-2.0mu\rho}$ is the real part of the frequency and $\tilde{\xi}_{\bm{q}}$ is the same thermal noise as in Eq. (20). There is no elastic restoring force in Eq. (22), which is analogous to the equation of motion for a colloidal particle in a fluid. Just as a colloid diffuses freely at long times [60], the variance of membrane height undulations grows linearly in time according to Eq. (22). Eventually, the membrane height will grow to a point where our assumption of linearity is no longer valid. At this point, membrane bending and surface tension forces will presumably prevent height fluctuations from continuing to grow—yet a description of such behavior is beyond the scope of the present work. Thus, irrespective of whether our ansatz is valid, membrane fluctuations are not expected to satisfy the equipartition result outside the dome where $q\notin(q_{0}^{-},q_{0}^{+})$.
Figure 3: Analysis of GUV contour fluctuations, modified due to membrane semipermeability. (a) Canonical protocol to calculate the bending modulus and surface tension of a fluctuating GUV. In a spherical vesicle of radius $r_{\mkern-1.0mu\mkern-1.0mu\text{v}}$ (left), the equatorial cross-section is imaged (center) at successive time intervals. Radial deviations from a circle are recorded as $h(x,t)$, where $x$ is the arclength along the unperturbed circle. The shape disturbances $h(x,t)$ are decomposed into one-dimensional Fourier modes, whose thermally-averaged amplitudes (right; adapted from Ref. [53]) are fit to Eq. (23) by tuning the values of $k_{\text{b}}$ and $\lambda_{\text{c}}$ (right; solid orange line) [29]. Theory and experiment deviate at large $q_{x}$ due to finite optical resolution in the latter. (b) Limitations of the equipartition result (21) due to solutes in the surrounding fluid. (left) Modes inside the dome (red lines) satisfy Eq. (21); those outside the dome (green dots) do not. Membrane undulations cannot be visualized above $q_{\text{opt}}\approx 0.025$ nm-1 (gray shading) due to limitations in the optical resolution [31]. (right) For a GUV with surface tension $\lambda_{\text{c}}\in(0,\lambda_{\text{c}}^{\mkern-1.0mu\ast})$, modes are characterized in the $q_{x}$$q_{y}$ plane. The equipartition result (21) is limited to modes with amplitude $q>q_{0}^{-}$, and large-$q$ modes cannot be visualized. The domains of integration in Eq. (23), corresponding to two different values of $q_{x}$, are shown as dashed vertical lines. The result of Eq. (23) is invalid when $q_{x}<q_{0}^{-}$, as in such cases the integral contains modes outside the dome. (c) Plot of experimental membrane fluctuations, over a range of reported surface tensions $\lambda_{\text{c}}$, from prior studies: Faizi et al. [57] ($3.1\!\cdot\!10^{-6}$ pN//nm), Rautu et al. [56] ($1.4\!\cdot\!10^{-5}$ pN//nm), Pécréaux et al. [29] ($1.7\!\cdot\!10^{-4}$ pN//nm), Park et al. [55] ($7\!\cdot\!10^{-4}$$3\!\cdot\!10^{-3}$ pN//nm), and Takatori & Sahu [53] ($4\!\cdot\!10^{-3}$ pN//nm). The $y$-axis is chosen to be independent of the vesicle radius, and depends only on $\lambda_{\text{c}}$ and $k_{\text{b}}$ [see Eq. (23)]. At large $q_{x}$, the data saturates due to optical resolution limitations. All data points below the dashed black line are outside the dome, and should be excluded when determining membrane parameters. Fluctuation data from the high-tension vesicle in Ref. [54] is not publicly available, and so cannot be plotted—though the first 88 modes should be excluded based on the reported value of $\lambda_{\text{c}}=0.025$ pN//nm [25, §4].
The standard contour analysis of a GUV. Figure 3(a) portrays the canonical analysis of a GUV undergoing thermal fluctuations, the details of which can be found in Ref. [29]. In short, the equatorial cross-section of a GUV is imaged at frequent time intervals. For each snapshot, the radial deviation from a circle of radius $r_{\mkern-1.0mu\mkern-1.0mu\text{v}}$ is measured and decomposed into normal modes. As GUVs are large and weakly curved, it is often convenient to apply the planar membrane theory [29]—which is simpler than its spherical counterpart [56, 57, 58, 31]. The position $x$ is defined as the distance along the circle, with $h(x,t)$ the radial perturbation, $\ell_{\text{c}}=2\mkern 1.0mu\pi\mkern 1.0mur_{\mkern-1.0mu\mkern-1.0mu\text{v}}$, and $q_{x}=2\mkern 1.0mu\pi\mkern 1.0mum/\ell_{\text{c}}=m/r_{\mkern-1.0mu\mkern-1.0mu\text{v}}$ for positive integers $m$. The one-dimensional (1D) fluctuation spectrum $\langle\lvert\mkern 1.0mu\tilde{h}_{q_{x}}\mkern-1.0mu\rvert^{2}\rangle$ is calculated from experimental measurements as in Fig. 3(a). To compare the 1D data with the theoretical 2D equipartition result of Eq. (21), the latter is averaged over all $q_{y}$ as [29]
[…]
The bending modulus $k_{\text{b}}$ and surface tension $\lambda_{\text{c}}$ are chosen so as to minimize the difference between the measured fluctuation spectrum and prediction of Eq. (23)—which, in Fig. 3(a), are respectively shown as black circles and an orange line. Note that short-wavelength fluctuations cannot be visualized in experiments due to a lower bound $\ell_{\text{opt}}\approx 250$ nm on the optical resolution, which sets an upper limit $q_{\text{opt}}:=2\mkern 1.0mu\pi/\ell_{\text{opt}}\approx 0.025$ nm-1 beyond which undulations cannot be observed [31]. In practice, short-wavelength modes are excluded when fitting $k_{\text{b}}$ and $\lambda_{\text{c}}$. Techniques also exist to correct for the finite camera integration time [29] and vertical focal projections [56], but are not discussed here.
The modified contour analysis of a GUV. When solutes are present in the fluid surrounding the membrane, the 2D equipartition result (21) is only valid for modes $\bm{q}$ with magnitude $\lvert\bm{q}\rvert\in(q_{0}^{-},q_{0}^{+})$. We seek to determine how the experimental analysis of contour fluctuations is altered by such a restriction. To this end, we first note that $q_{\text{opt}}<q_{0}^{+}$, as shown in Fig. 3(b). Short-wavelength undulations outside the dome thus cannot be observed in experiments, and do not contribute to the determination of $k_{\text{b}}$ and $\lambda_{\text{c}}$. Long-wavelength fluctuations outside the dome, on the other hand, are captured experimentally. Figure 3(b) characterizes modes in the $q_{x}$$q_{y}$ plane as being either outside the dome ($\lvert\bm{q}\rvert<q_{0}^{-}$), inside the dome and visually accessible ($q_{0}^{-}<\lvert\bm{q}\rvert<q_{\text{opt}}$), or visually inaccessible ($\lvert\bm{q}\rvert>q_{\text{opt}}$). Here, domains of the contour integral in Eq. (23) for two values of $q_{x}$ are shown as dashed vertical lines. If $q_{x}<q_{0}^{-}$, this domain contains modes for which the equipartition result (21) does not apply. The result of Eq. (23), which relies on Eq. (21), is accordingly valid only when $q_{x}>q_{0}^{-}$—with the understanding that as $q_{x}$ increases, fewer modes are optically accessible.
To determine $k_{\text{b}}$ and $\lambda_{\text{c}}$ from GUV contour fluctuations, it is useful to plot fluctuation amplitudes as in Fig. 3(c)—which contains experimental data from several prior investigations [57, 56, 29, 55, 53]. Here, both the $x$-axis and $y$-axis are chosen to be independent of $r_{\mkern-1.0mu\mkern-1.0mu\text{v}}$ [cf. Eq. (23)]. Next, experimental data inside and outside the dome need to be identified. To this end, we combine Eqs. (10) and (23) to obtain the fluctuation amplitudes of 1D modes $q_{x}=q_{0}^{-}$ on the dome, for tensions $\lambda_{\text{c}}\in(0,\lambda_{\text{c}}^{\mkern-1.0mu\ast})$:
[…]
Importantly, $k_{\text{b}}$ is not known a priori, and so Eq. (24) can be approximated as $\ell_{\text{c}}\mkern 1.0mu\langle\lvert\mkern 1.0mu\tilde{h}_{q_{x}}\mkern-1.0mu\rvert^{2}\rangle\approx k_{\text{B}}\theta/(8\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq_{x}{}^{\!\!2}D)$—valid when $k_{\text{b}}\mkern 1.0muq_{x}\ll 8\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muD$—in a first analysis of the data. Equation (24) is plotted as the dashed black line in Fig. 3(c); all of the data points below it are outside the dome and should be excluded when determining $k_{\text{b}}$ and $\lambda_{\text{c}}$.
Our aggregation and analysis of the experimental data in Fig. 3(c) is detailed in §4 of the SM [25]. Most often, vesicle tensions $\lambda_{\text{c}}\!\sim\!10^{-7}$$10^{-4}$ pN//nm are low [29, 56, 57, 58] and all data lie inside the dome. As it turns out, a vesicle at low tension is experimentally favorable because its undulations are large and hence easy to image. In many cases, the vesicle’s osmotic environment is manipulated to elicit a low tension. Notable exceptions are the so-called active vesicles containing either self-propelled Janus particles or motile bacteria [53, 54, 55]. For these systems, moderate vesicle tensions are required for active particles to not overcome the elastic membrane restoring force and form a tube. Given the fluctuation data for active vesicles in Fig. 1 of Ref. [53] and Fig. 6(c) of Ref. [55], between one and eight long-wavelength modes lie outside the dome [25, §4]—though the overall findings of these studies are unaffected. In contrast, for the high-tension ($\lambda_{\text{c}}=0.025$ pN//nm) active vesicle in Fig. 1(g) of Ref. [54], the first 88 long-wavelength modes lie outside the dome. It is likely that upon excluding these modes, the calculated values of $k_{\text{b}}$ and $\lambda_{\text{c}}$ would be affected. Unfortunately, the experimental fluctuation data for this vesicle is not provided, and we cannot investigate further. Nevertheless, we find osmotic forces can indeed alter membrane fluctuations in biologically-relevant scenarios.
Concluding remarks
In this work, we investigated the linearized dynamics of a planar lipid membrane, surrounded by a Newtonian fluid in which solutes are dissolved. Though the lipid bilayer is only weakly permeable to water, the presence of solutes can significantly alter its dynamics. More specifically, when $\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}<\tilde{\omega}_{\mkern-1.0mu\text{m}}$ and diffusion is slow relative to the impermeable membrane, undulations of the semipermeable system no longer decay at a frequency near $\tilde{\omega}_{\mkern-1.0mu\text{m}}$. In such cases, either ($i$) the ansatz that all quantities can be expressed in terms of planar normal modes as in Eqs. (2), (3), and (12) is no longer valid, or ($i\mkern-1.0mui$) the bilayer only decays at the fast inertial frequency $\omega_{\mkern-2.0mu\rho}\gg\omega_{\mkern-1.0mu\text{m}}$. When thermal perturbations are incorporated into the theory, the well-known equipartition result (21) is found to only be experimentally relevant when $q\in(q_{0}^{-},q_{\text{opt}})$—a range that narrows as $\lambda_{\text{c}}$ is increased. Our results are relevant when membrane tensions are large, as is the case with active vesicles [53, 54, 55] and a variety of biological scenarios [61, 30, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72].
The vanishing of the membrane mode at low $q$ and large $\lambda_{\text{c}}$ opens several lines of inquiry. We are particularly interested to see if membranes in spherical or cylindrical geometries respond in the same fashion; the former is needed for a careful analysis of GUV undulations [56, 57, 58] while the latter could be relevant to neuronal systems [61, 30, 63, 62, 64]. In all scenarios where the membrane branch vanishes, the behavior of the coupled membrane–fluid–solute system can be interrogated with nonlinear simulations [46, 47, 48, 49, 50, 51, 52]. Additional levels of complexity can be introduced by considering membranes that are not ideally selective [3] or charged solutes—with the latter requiring extensive theoretical developments [73, 74, 75, 76, 77].
Data, methods, and software availability. The Supplemental Material [25] details the theoretical techniques used, as well as the methods of numerical solution. The code used to generate all data in the present study is publicly available at github.com/sahu-lab/osmosis-flat.
Acknowledgments. We express our most sincere gratitude to Prof. Kranthi Mandadapu for his perspective on lipid membrane behavior. We are also thankful to have discussed Ref. [3] with Dr. Ahmad Alkadri and Kranthi, as well as Refs. [59, 55] with Dr. Kisung Lee, Dr. Myoenggon Park, and Prof. Steve Granick. It is a pleasure to thank Dr. Joël Tchoufag for many stimulating discussions, Profs. Brian Camley and Tine Curk regarding their thoughts on membrane fluctuations, and Prof. Sho Takatori for introducing us to the experimental analysis of GUV fluctuations.